Please check the answers below the table and help with a question (1-4). Thanks   Name Show your calculations for moles of lz reacted: (find through moles of S,0,?) Show your calculation for molarity of la r1. What is the overall order of the reaction in this lab? 2. Use your Rate Law and average k to calculate the reaction rate w
Answer #1

1) The overall order is 2.

2) The rate is calculated:

Rate = 5.23x10 ^ -4 * (0.025) * (0.032) = 4.18x10 ^ -7 M / s

3) a) Write the rate law:

Rate = k * [A] * [B] ^ 2

b) k has units of: M-2 * s-1

c) The reaction rate will be doubled.

4) It has the integrated expression for second order:

1 / [X] = k * t + 1 / [X] o

You have:

[X] = 1 / [k * t + 1 / [X] o] = 1 / [5.5x10 ^ -2 * 89 + 1 /2.5] = 0.189 M

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.